40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 7 14:04:48 2011, steve, Update, SUS, He/Ne intensity noise effect on oplevs 7x
    Reply  Fri Oct 7 23:23:05 2011, kiwamu, Update, SUS, SRM oplev was oscillating 
       Reply  Wed Oct 19 10:49:35 2011, kiwamu, Update, SUS, SRM oscillation removed SRMoplevKWMtune.png
    Reply  Wed Oct 12 08:43:30 2011, steve, Update, SUS, ITMX oplev improved a bit ITMXoplev.pngITMXoplevservo_ON.png
       Reply  Thu Oct 13 14:23:09 2011, steve, Update, SUS, ITMX oplev with 3 mm beam on qpd ITMXoplev1103p#2.pngITMXoplev.png
          Reply  Mon Oct 17 17:07:30 2011, steve, Update, SUS, ETMX oplev returning beam od 3 mm ETMX-oplev.jpg
             Reply  Wed Oct 19 09:44:10 2011, steve, Update, SUS, ETMX oplev power spectrum ETMXoplev2lens.png
             Reply  Fri May 18 15:25:15 2012, steve, Update, SUS, AR coated lens swapped in at ETMX oplev ETMXoplevlensAR.png
Message ID: 5655     Entry time: Wed Oct 12 08:43:30 2011     In reply to: 5630     Reply to this: 5660
Author: steve 
Type: Update 
Category: SUS 
Subject: ITMX oplev improved a bit 

 Atm2 is before optical path adjustment. The idea was to remove possible clipping in vacuum.

Coherense significantly reduced below 4 Hz

Today I will replace the He/Ne laser 1125P with 1103P

 

Attachment 1: ITMXoplev.png  53 kB  | Hide | Hide all
ITMXoplev.png
Attachment 2: ITMXoplevservo_ON.png  52 kB  | Hide | Hide all
ITMXoplevservo_ON.png
ELOG V3.1.3-